# RAGtifier: Evaluating RAG Generation Approaches of State-of-the-Art RAG Systems for the SIGIR LiveRAG Competition

### v1

@@ Abstract.
Retrieval-Augmented Generation (RAG) enriches Large Language Models (LLMs) by combining their internal, parametric knowledge with external, non-parametric sources, with the goal of improving factual correctness and minimizing hallucinations. The LiveRAG 2025 challenge explores RAG solutions to maximize accuracy on DataMorgana’s QA pairs, which are composed of single-hop and multi-hop questions. The challenge provides access to sparse OpenSearch and dense Pinecone indices of the Fineweb 10BT dataset. It restricts model use to LLMs with up to 10B parameters and final answer generation with Falcon-3-10B. A judge-LLM assesses the submitted answers along with human evaluators. By exploring distinct retriever combinations and RAG solutions under the challenge conditions, our final solution emerged using InstructRAG in combination with a Pinecone retriever and a BGE reranker. Our solution achieved a correctness score of $1.13$ and a faithfulness score of $0.55$, placing fourth in the SIGIR 2025 LiveRAG Challenge [^1] Non-final LLM scoring rankings from the organizer. [Accessed 22-05-2025]. The RAGtifier code is publicly available [^2] https://git.l3s.uni-hannover.de/liverag/ragtifier.
@@ 4.3. Retriever Performance
Considering the time constraints of the LiveRAG Challenge and the influence of golden documents in retrieval@k, we investigate the runtime of the retrieval and the number of golden documents returned at distinct retrieval@k values for the provided OpenSearch and Pinecone indices. Figure 4 shows that the number of gold documents increases continuously with higher retrieval@k. Notably, Pinecone outperforms OpenSearch at @$k=20$ for multi-hop questions and at @$k=50$ for single-hop questions. Runtime measurements (Figure 5) reveal that Pinecone is faster for all retrieval@k at $k=[1.600]$. OpenSearch takes $0.12s$ and Pinecone $0.15s$ at @$k=1$, scaling nearly linearly up to $k=600$, where OpenSearch takes $0.9s$ and Pinecone $0.58s$. We decide to proceed with Pinecone due to its performance after reaching Falcon-3-10B’s context limit of about $50$ retrieved documents. Consequently, we check the performance increase by using a reranker in combination with retriever@$k>50$.
@@ 4.4. Reranker Performance
The runtime of BGE (Figure 5) increases with the number of retrieved documents. Further experiments reveal that BGE takes $\sim 11.2s$ to rerank 400 retrieved documents. Considering the LiveRAG time constraints and available computational resources, we limit further experiments to 300 retrieved documents, which takes $\sim 8.6s$ per reranking operation per question. If we increase k up to the context limit of $k=50$ (Figure 4), the percentage of returned gold documents increases due to the query alone. We hypothesize an increasing RAG performance due to more golden documents are present in the context when a higher retrieval@k is used in combination with a reranker set to $k\leq 20$. Therefore, we searched within the subset of viable retriever and reranker settings for a configuration that outperforms retrieval@50 in terms of gold documents@k. For single-hop questions using Pinecone@300 with BGE@10 and OpenSearch and Pinecone@[100, 300] with BGE@20, more gold documents remaining in the reranked set compared to using OpenSearch or Pinecone@50 alone. Similarly, for multi-hop questions, this occurs for OpenSearch and Pinecone@[100, 300] with BGE@20.
@@ 4.5. Final System Performance
With our insights into retriever and reranker performance, we cross-evaluate various settings. We use Pinecone@$\{100,200,300\}$ with BGE@$\{5,8,10,12\}$ combined with optional inverted context order for each RAG solution. Due to the iterative context generation of IterDRAG, we omitted the context ordering and tested Pinecone@$\{100,200\}$ with BGE@$\{5,10\}$ for the initial retrieval step. For additional retrieval steps in IterDRAG, we considered Pinecone@$200$ with BGE@$\{4,5\}$ for a maximum of four and five iterations respectively, and BGE@3 for six iterations.
Considering the LiveRAG constraints, increasing retrieval@k for a fixed rerank@k does not consistently lead to better performance. The performance differences due to a change in retrieval@k remain mainly within a ±$\pm 2\%$ range of the correctness evaluation metric (Figure 2). Increasing rerank@k with fixed retrieval@k results in a higher variation in the correctness score, where we measure variations from $1\%$ up to $25\%$. The influence of the inverted context order averages to $1\%$ performance increase.
The best-performing RAG approaches are listed in Figure 1, using identical settings of Pinecone@200, BGE@5 and inverted context order. IterDRAG uses Pinecone@200 and BGE@10 for initial retrieval, and Pinecone@200 with BGE@4 for up to 5 iterations. With InstructRAG and IterDRAG perform comparably on Gemma-3-27B, we select both as possible approaches for LiveRAG. During the live challenge day, we generated answers with both RAG solutions and evaluated the results using Gemma-3-27B, omitting the golden document and golden answer, supplemented by human evaluation. Resulting in InstructRAG outperforms IterDRAG in terms of the evaluation metric. IterDRAG achieved an average correctness of $1.70$ and a faithfulness score of $0.73$, while InstructRAG achieved a correctness of $1.91$ ($+28.1\%$) and a faithfulness score of $0.93$ ($+27.4\%$). An additional manual comparison between these two approaches revealed an occasionally subjectively better question-answer alignment for InstructRAG. Compared to our measurements, the organizer’s LLM evaluation returned lower scores for our submitted InstructRAG-based approach: correctness of $1.13$ ($-40.9\%$) and faithfulness of $0.55$ ($-40.9\%$). This difference was likely influenced by using a more capable judge LLM and accessing the golden document and golden answer.

### v2

@@ Abstract.
Retrieval-Augmented Generation (RAG) enriches Large Language Models (LLMs) by combining their internal, parametric knowledge with external, non-parametric sources, with the goal of improving factual correctness and minimizing hallucinations. The LiveRAG 2025 challenge explores RAG solutions to maximize accuracy on DataMorgana’s QA pairs, which are composed of single-hop and multi-hop questions. The challenge provides access to sparse OpenSearch and dense Pinecone indices of the Fineweb 10BT dataset. It restricts model use to LLMs with up to 10B parameters and final answer generation with Falcon-3-10B. A judge-LLM assesses the submitted answers along with human evaluators. By exploring distinct retriever combinations and RAG solutions under the challenge conditions, our final solution emerged using InstructRAG in combination with a Pinecone retriever and a BGE reranker. Our solution achieved a correctness score of $1.13$ and a faithfulness score of $0.55$ in the non-human evaluation, placing it overall in third place in the SIGIR 2025 LiveRAG Challenge (Carmel et al., 2025). The RAGtifier code is publicly available [^1] https://git.l3s.uni-hannover.de/liverag/ragtifier.
@@ 4.3. Retriever Performance
Considering the time constraints of the LiveRAG Challenge and the influence of golden documents in retrieval@k, we investigate the runtime of the retrieval and the number of golden documents returned at distinct retrieval@k values for the provided OpenSearch and Pinecone indices. Figure 2 shows that the number of gold documents increases continuously with higher retrieval@k. Notably, Pinecone outperforms OpenSearch at @k=20@k=20 for multi-hop questions and at @k=50@k=50 for single-hop questions. Runtime measurements (Figure 4) reveal that Pinecone is faster for all retrieval@k at $k=[1.600]$. OpenSearch takes $0.12s$ and Pinecone $0.15s$ at @k=1@k=1, scaling nearly linearly up to $k=600$, where OpenSearch takes $0.9s$ and Pinecone $0.58s$. We decide to proceed with Pinecone due to its performance after reaching Falcon-3-10B’s context limit of about $50$ retrieved documents. Consequently, we check the performance increase by using a reranker in combination with retriever@$k>50$.
@@ 4.4. Reranker Performance
The runtime of BGE (Figure 4) increases with the number of retrieved documents. Further experiments reveal that BGE takes $\sim 11.2s$ to rerank 400 retrieved documents. Considering the LiveRAG time constraints and available computational resources, we limit further experiments to 300 retrieved documents, which takes $\sim 8.6s$ per reranking operation per question. If we increase k up to the context limit of $k=50$ (Figure 2), the percentage of returned gold documents increases due to the query alone. We hypothesize an increasing RAG performance due to more golden documents are present in the context when a higher retrieval@k is used in combination with a reranker set to $k\leq 20$. Therefore, we searched within the subset of viable retriever and reranker settings for a configuration that outperforms retrieval@50 in terms of gold documents@k (Figure 3). For single-hop questions using Pinecone@300 with BGE@10 and OpenSearch and Pinecone@[100, 300] with BGE@20, more gold documents remaining in the reranked set compared to using OpenSearch or Pinecone@50 alone. Similarly, for multi-hop questions, this occurs for OpenSearch and Pinecone@[100, 300] with BGE@20.
@@ 4.5. Final System Performance
[…]
Considering the LiveRAG constraints, increasing retrieval@k for a fixed rerank@k does not consistently lead to better performance. The performance differences due to a change in retrieval@k remain mainly within a $\pm 2\%$ range of the correctness evaluation metric (Figure 5). Increasing rerank@k with fixed retrieval@k results in a higher variation in the correctness score, where we measure variations from $1\%$ up to $25\%$. The influence of the inverted context order averages to $1\%$ performance increase.
Table 1 presents the results of the best-performing RAG settings for each approach, using identical settings of Pinecone@200, BGE@5 and inverted context order. IterDRAG uses Pinecone@200 and BGE@10 for initial retrieval, and Pinecone@200 with BGE@4 for up to 5 iterations. With InstructRAG and IterDRAG perform comparably on Gemma-3-27B, we select both as possible approaches for LiveRAG. During the live challenge day, we generated answers with both RAG solutions and evaluated the results using Gemma-3-27B, omitting the golden document and golden answer, supplemented by human evaluation. Resulting in InstructRAG outperforms IterDRAG in terms of the evaluation metric. IterDRAG achieved an average correctness of $1.70$ and a faithfulness score of $0.73$, while InstructRAG achieved a correctness of $1.91$ ($28.1\%$) and a faithfulness score of $0.93$ ($27.4\%$). An additional manual comparison between these two approaches revealed an occasionally subjectively better question-answer alignment for InstructRAG. Compared to our measurements, the organizer’s LLM evaluation returned lower scores for our submitted InstructRAG-based approach: correctness of $1.13$ ($40.9\%$) and faithfulness of $0.55$ ($40.9\%$). This difference was likely influenced by using a more capable judge LLM and accessing the golden document and golden answer.
@@ Appendix A Appendix
Figure 3. Gold documents [%] in retrieval for Faclon3-10B context window of k=50 documents with Pinecone@k (76.2%) compared to a selection of Pinecone@k1 and BGE@k2. Pinecone@80 with BGE@20 (76.7%) reduces the context size while exceeding Pinecone@50 performance.